OpenStudy (anonymous):

6x ^{3}-24x ^{2}div x ^{2}-3x+2 * x ^{2} +8x+16div 2x ^{2}+12x

jhonyy9 (jhonyy9):

so if you want writing this : 6x*3 - 24x*2 x*2 +8x +16 6x*2(x-4)(x+4)*2 ------------ x ------------- = ---------------- = x*2 -3x +2 2x*2 +12x (x-1)(x-2)2x(x+6) 3x(x-4)(x+4)*2 = -------------- (x-1)(x-2)(x+6)

OpenStudy (anonymous):